# Pietro Cassieri+, Gianluca Fimiani+, Simone Romano+, Marina Garofano+, Alessia Bramanti+, Giuseppe Scanniello+

### v1

@@ V Result
We report the most important preliminary results achieved in the early stage of PrediHealt arranged according to O1, O2, and O3 (i.e., the objectives for which we obtained results).
@@ V-A O1 - Interoperable Web-Based IoT Platform
The selection process prioritized platforms that seamlessly integrate with wearable devices and environmental sensors while ensuring compliance with evolving regulations and technological advancements. Platforms that lacked active maintenance or regular updates were excluded, as outdated solutions pose significant risks in terms of security, compatibility with modern devices, and adaptability to AI-driven predictive models. In contrast, modern and actively supported platforms with strong IoT integration were deemed more suitable for the project’s long-term sustainability. Further details on the selection process are deliberately omitted due to space constraints and as they fall outside the scope of this paper. In Table III, we report the final set of platforms from which we selected the one for PrediHealth, i.e., FIWARE [25].
FIWARE is a web-based IoT platform supported by the European Union and released under a permissive open-source license (BSD License). Backed by an active open-source community, FIWARE remains up-to-date and widely adopted in healthcare, with proven applications in hospital automation and real-time patient monitoring. In addition, its compliance with European digital health regulations provides strategic advantages, including potential funding opportunities. The modular architecture, based on Generic Enablers, allows customized remote health monitoring solutions. Scalability is key, as PrediHealth requires real-time processing of growing IoT data while maintaining high performance. FIWARE’s compatibility with Docker and Kubernetes enhances deployment flexibility across cloud and edge environments. This platform, through to the Context Broker, efficiently manages real-time data from IoT sensors and wearables, ensuring accurate patient monitoring and informed decision-making. FIWARE also integrates advanced data processing tools for distributed computing, big data analytics, and ML, enabling AI-driven predictive healthcare. FIWARE is designed with a strong focus on security, incorporating multiple layers of protection to ensure data integrity, confidentiality, and controlled access. It implements robust authentication mechanisms through OAuth2, enabling secure authorization and identity verification for users and applications. Additionally, XACML-based access control policies allow for fine-grained permission management, ensuring that only authorized entities can access specific resources based on predefined rules. These features let FIWARE be compliant with community security standards and regulations. FIWARE also supports diverse IoT protocols, including MQTT, LoRaWAN, SigFox, OPC-UA, and Lightweight M2M, ensuring seamless device integration and interoperability and compliance with healthcare standards. We designed and implemented an API to expose healthcare data in the HL7 FHIR format.
@@ V-B O2 - Telemonitoring Kit for HF Patients
The design of the telemonitoring kit for HF patients is based on the integration of environmental sensors and advanced (wearable and not) medical devices. As for the environmental sensors, monitoring environmental parameters is essential for early detection of cardiovascular deterioration because parameters including air quality, temperature, and humidity complement the core vital signs in assessing cardiovascular health. Air quality monitoring was included due to the established link between air pollution and cardiovascular risks. For example, Khraishah et al. [26] showed that exposure to pollutants contributes to metabolic and cardiovascular diseases and that wildfire-related air pollution exacerbates cardiovascular conditions. Similarly, temperature and humidity monitoring were prioritized given their impact on cardiovascular mortality, as extreme heat has been linked to increased cardiovascular-related deaths [27]. We also considered environmental noise, given its association with hypertension and cardiovascular disease [28, 29, 30, 31]. We conducted a study on various environmental sensors. The results indicated that, although there were differences in the characteristics of the sensors, none of them clearly outperformed the others. Consequently, the selection process prioritized factors such as cost-effectiveness, measurement accuracy, and adherence to European manufacturing standards. This approach aligns with the requirements for projects funded by the National Recovery and Resilience Plan (PNRR), which suggests that the hardware used must be compliant with European standards/certifications. Table IV shows the selected sensors and their technical specifications.
@@
Also, continuous weight monitoring is crucial for HF patients, as fluid retention can indicate worsening HF, as well as physical activity, as regular exercise has been shown to prevent and manage cardiovascular disease [38, 39]. Data gathered with medical devices in PrediHealth include body temperature, heart rate, HRV, oxygen saturation (SpO2), blood pressure (both systolic and diastolic), and ECG data. These parameters were chosen because they serve as critical indicators for detecting early signs of hemodynamic instability, autonomic dysfunction, or respiratory distress—key factors in managing HF [40, 41, 42, 43]. More specifically, body temperature is crucial for identifying potential infections or inflammatory processes, which could signal a deterioration in the patient’s condition. Heart rate is a cardiovascular indicator that, in case of any abnormal increase or decrease, points to potential cardiac stress or arrhythmias. HRV, which reflects autonomic function, is particularly valuable for assessing the balance between the sympathetic and parasympathetic nervous systems, with low HRV being associated with an increased risk of cardiovascular events. SpO2 is an important marker of respiratory function, with low levels indicating potential respiratory failure or pulmonary edema, which is common in HF patients. Blood pressure, both systolic and diastolic, is important for monitoring hypertension, a significant risk factor for HF. Fluctuations in blood pressure may indicate a worsening cardiac function or the development of pulmonary hypertension. Additionally, ECG data provide detailed insights into the heart’s electrical activity, allowing for the detection of arrhythmias, ischemia, and other abnormalities that may require immediate intervention. The integration of these gathered data with physiological information supports the early detection of potential health deterioration and allows for proactive interventions.
@@
We focused on devices that ensure high measurement accuracy, seamless integration with our IoT platform, and compliance with medical standards (e.g., HL7 FHIR). A comprehensive analysis of commercially available devices was conducted, encompassing smartwatches, smartbands, sensor-integrated smart garments, and weight-monitoring devices, all of which play a crucial role in HF telemonitoring. We considered scientific literature and gray literature to ensure a thorough assessment of available technologies. The selected devices were assessed based on multiple criteria, including their ability to provide real-time physiological and behavioral data, interoperability with third-party healthcare applications, and accessibility through APIs or SDKs. Due to space constraints, Table V shows only the selected devices. Specifically, we selected the Withings ScanWatch 2 as the primary wearable device. This choice was driven by its ability to record a single-lead ECG directly from the wrist, enabling early detection of arrhythmias such as atrial fibrillation. This device also provides continuous monitoring of heart rate, HRV, SpO2, and blood pressure, offering a comprehensive cardiovascular assessment. A key advantage of ScanWatch 2 is its seamless integration with IoT platforms via Withings’ well-documented APIs. These APIs ensure secure access to real-time health data (e.g., ECG traces), facilitating efficient data collection and analysis. Moreover, ScanWatch 2 meets stringent medical certification standards, ensuring measurement accuracy and compliance with regulatory requirements. Compared to other smartwatches, ScanWatch 2 offers a balance between advanced medical-grade features and ease of integration, making it an optimal solution for telemonitoring. For similar reasons, we selected the Withings Body+ scale.
@@
Wearable Devices
Withings ScanWatch 2 - ECG - Heart Rate - Temperature Variation - SpO2 - Physical Activity - Sleep Monitoring Yes 30
Scales
Withings Body+ - Weight - BMI - Fat mass - Muscle mass - Bone mass - Water percentage Yes 540
@@ V-C O3 - Developing tools to support decision-making
We implemented the predictive model for stratification using clinical and echocardiographic features extracted from a dataset of 182 elements. We report the used features in Table VI. The model architecture is based on an ensemble learning technique, specifically a variation of stacking [44]. In standard stacking, multiple base models (often different from each other) are trained on the same set of features, and a meta-model then combines their outputs to generate the final prediction. However, in our approach, instead of using different machine learning models on the same feature set, we employed two identical models trained separately: one using only clinical features and the other using only echocardiographic features. The outputs of these two models serve as inputs for the meta-model, which provides the final prediction shown to the physician. The output of this metamodel is expressed in terms of estimated lifespan. However, due to the sensitivity of the medical domain and the precision requirements of clinical data, patients for whom the model predicts a lifespan of fewer than 440 days are classified as being at cardiovascular risk.
The performance of our predictive model was assessed using metrics that are commonly employed in the medical domain. Specifically, we considered: (i) Accuracy ($\frac{TP+TN}{TP+TN+FP+FN}$), (ii) Precision ($\frac{TP}{TP+FP}$), (iii) Sensitivity/Recall ($\frac{TP}{TP+FN}$), (iv) F1-Score ($2\times\frac{\text{Precision}\times\text{Sensitivity}}{\text{Precision}+\text{Sensitivity}}$), and (v) Diagnostic Odds Ratio ($\frac{TP\times TN}{FP\times FN}$). For the sake of clarity, TPs (True Positives) refer to patients correctly classified as people at risk (with lifespan less than 440 days),while TNs (True Negatives) refer to patients correctly identified as people with a lifespan greater than 440 days. Instead, FPs (False Positives) represent patients mistakenly classified as at risk, while FNs (False Negatives) represent patients mistakenly identified as at risk.
@@
Metric Value Description
Accuracy 0.84 84% of cases are correctly classified.
Precision (At-Risk) 0.79 79% of patients flagged as at risk truly are at risk.
Sensitivity (At-Risk) 0.95 95% of at-risk patients are correctly identified.
F1-Score (At-Risk) 0.86 Good balance between precision and sensitivity.
Diagnostic Odds Ratio 45.6 Strong discriminative ability.
The results obtained from the model are summarized in Table VII. These results suggest that the model performs pretty well. Although an accuracy of 84% can be considered moderate in some ML contexts, it is acceptable in the PrediHealth context given the priority of capturing high-risk cases. The high sensitivity (95%) ensures that nearly all high-risk patients are detected, while a relatively low precision (79%) resulting from a high number of false positives is an acceptable trade-off in our project. The good balance between sensitivity and precision is shown by the obtained F1-Score value (86%). The high Diagnostic Odds Ratio (45.6) confirms the model’s robust ability to discriminate between high-risk and low-risk patients.
As for the thresholds/scores for risk factors for HF exacerbations when using telemonitoring kits, we chose them on clinical evidence and guidlines [45, 46, 47]. Specifically, threshold values considered at risk include SpO2 $<$ 92%, resting heart rate (HR) $>$ 100 bpm (tachycardia) or $<$ 50 bpm (bradycardia), and persistently low heart rate variability (HRV), typically SDNN $<$ 20 ms. Early signs of decompensation can be inferred when body weight increases by more than 2 kg in three days, suggesting fluid overload. Other critical indicators include systolic blood pressure $>$ 140 mmHg or $<$ 90 mmHg, and diastolic pressure $>$ 90 mmHg or $<$ 60 mmHg, which may indicate hemodynamic imbalance. Finally, a respiratory rate $>$ 20 breaths/min and body temperature $>$ 37.5 or $<$ 36 may signal infection, hypoperfusion, or cardiovascular stress. Finally, a single-lead ECG, validated for atrial fibrillation detection by identifying irregular RR intervals and absent P waves, enables early recognition of AF, which is critical as it can impair cardiac output and increase the risk of decompensation and thromboembolic events. When risk factors occur, the IoT platform will notify both the patient and the appropriate healthcare infrastructure. This phase of PrediHealth is still ongoing. We aim to collect field data to further enhance the body of knowledge in this respect.
@@ VI Final Remarks and Expected Impact
PrediHealth has the potential to revolutionize the management of cardiovascular diseases, particularly Heart Failure (HF), by integrating state-of-the-art technologies such as telemonitoring, m-health, and AI/ML. It is centered around the development of a web-based IoT application designed for remote HF patient monitoring. This innovative solution enables the continuous monitoring of vital signs and environmental parameters, thereby improving patient engagement and allowing for real-time oversight. By leveraging AI/ML-powered algorithms, we provide support to clinical decision-making (e.g., patient stratification), monitor disease, and proactively prevent exacerbations. With its holistic and data-driven approach to integrating clinical and environmental factors, PrediHealth has the potential to transform HF management, foster personalized care, and contribute to the long-term healthcare sustainability.
@@ Acknowledgment
This project has been financially supported by the European Union NEXTGenerationEU project and by the Italian Ministry of the University and Research MUR (M4C2) as part of the project “Tuscany Health Ecosystem” - THE - Spoke 10 - CUP - J13C22000420001.

### v2

@@ V Result
In this section, we report the most important preliminary results achieved in the early stage of PrediHealt arranged according to O1, O2, and O3 (i.e., the objectives for which we obtained some preliminary results).
@@ V-A O1 - Interoperable Web-Based IoT Platform
The selection process prioritized platforms that integrate with wearable devices and environmental sensors while ensuring compliance with evolving regulations and technological advancements. Platforms that lacked active maintenance or regular updates were excluded, as outdated solutions pose significant risks in terms of security, compatibility with modern devices, and adaptability to AI-driven predictive models. In contrast, modern and actively supported platforms with strong IoT integration were deemed more suitable for the project’s long-term sustainability. Further details on the selection process are omitted due to space constraints and as they fall outside the main goal of this paper (introducing PrediHealth and its ongoing results). In Table III, we report the final set of platforms from which we selected FIWARE [15].
FIWARE is a web-based IoT platform supported by the European Union [15, 26] and released under a permissive open-source license (BSD License). Backed by an active open-source community, FIWARE remains up-to-date. It is widely adopted in healthcare, with proven applications in hospital automation and real-time patient monitoring. In addition, FIWARE compliance with European digital health regulations provides strategic advantages [16], including potential funding opportunities. The modular architecture of FIWARE, based on Generic Enablers, allows customized remote health monitoring solutions. Scalability is key, as PrediHealth requires real-time processing of growing IoT data while maintaining high performance. FIWARE’s compatibility with Docker and Kubernetes supports high scalability and enhances deployment flexibility across cloud and edge environments. FIWARE, through the Context Broker, efficiently manages real-time data from IoT sensors and wearables. It also integrates advanced data processing tools for distributed computing, big data analytics, and ML, enabling AI-driven predictive healthcare. FIWARE is designed with a strong focus on security, incorporating multiple layers of protection to ensure data integrity, confidentiality, and controlled access. It implements robust authentication mechanisms through OAuth2, enabling secure authorization and identity verification for users and applications. Additionally, XACML-based access control policies allow for fine-grained permission management, ensuring that only authorized entities can access specific resources based on predefined rules. These features let FIWARE be compliant with community security standards and regulations. FIWARE also supports diverse IoT protocols, including MQTT, LoRaWAN, SigFox, OPC-UA, and Lightweight M2M, ensuring seamless device integration, interoperability, and compliance with healthcare standards. We designed and implemented an API to expose healthcare data in the HL7 FHIR format by leveraging the FIWARE’s HL7 Smart Data Model [27].
@@ V-B O2 - Telemonitoring Kit for HF Patients
The design and the development of the telemonitoring kit for HF patients is based on the integration of environmental sensors and advanced (wearable and not) medical devices. As for the environmental sensors, monitoring environmental parameters is essential because parameters including air quality, temperature, and humidity complement the core vital signs in assessing cardiovascular health (e.g., [28, 29]). In particular, air quality monitoring was included due to the established link between air pollution and cardiovascular risks. Khraishah et al. [28] showed that exposure to pollutants contributes to metabolic and CVDs and that wildfire-related air pollution exacerbates cardiovascular conditions. Similarly, temperature and humidity monitoring were prioritized given their impact on cardiovascular mortality, as extreme heat has been linked to increased cardiovascular-related deaths [29]. We also considered environmental noise, given its association with hypertension and CVDs [30, 31, 32, 33]. We conducted a study on various environmental sensors. The results indicated that, although there were differences in the characteristics of the sensors, none of them outperformed the others. Consequently, the selection process prioritized factors such as cost-effectiveness, measurement accuracy, and adherence to European manufacturing standards. This approach aligns with the requirements for projects funded by the National Recovery and Resilience Plan (PNRR), which suggests that the hardware used must be compliant with European standards/certifications. Table IV shows the selected sensors and their technical specifications.
@@
Also, continuous weight monitoring is crucial for HF patients, as fluid retention can indicate worsening HF, as well as physical activity, as regular exercise has been shown to prevent and manage CVDs [40, 41]. The kind of data gathered in PrediHealth with medical devices were chosen because they serve as critical indicators for detecting early signs of hemodynamic instability, autonomic dysfunction, or respiratory distress—key factors in managing HF [42, 43, 44, 45]. More specifically, body temperature is crucial for identifying potential infections or inflammatory processes, which could signal a deterioration in the patient’s condition. Heart rate is a cardiovascular indicator that, in case of any abnormal increase or decrease, points to potential cardiac stress or arrhythmias. HRV, which reflects autonomic function, is particularly valuable for assessing the balance between the sympathetic and parasympathetic nervous systems, with low HRV being associated with an increased risk of cardiovascular events. SpO2 is an important marker of respiratory function, with low levels indicating potential respiratory failure or pulmonary edema, which is common in HF patients. Blood pressure, both systolic and diastolic, is important for monitoring hypertension, a significant risk factor for HF. Fluctuations in blood pressure may indicate a worsening cardiac function or the development of pulmonary hypertension. Additionally, ECG data provide detailed insights into the heart’s electrical activity, allowing for the detection of arrhythmias, ischemia, and other abnormalities that may require immediate intervention. The integration of these kinds of data with physiological information supports the early detection of potential health deterioration and allows for proactive interventions.
@@
We focused on devices that ensure high measurement accuracy, seamless integration with our IoT platform, and compliance with medical standards (e.g., HL7 FHIR). A comprehensive analysis of commercially available devices was conducted, encompassing smartwatches, smartbands, sensor-integrated smart garments, and weight-monitoring devices, all of which could play a role in HF telemonitoring. We considered scientific literature and gray literature to ensure a thorough assessment of available technologies. The selected devices were assessed based on multiple criteria, including their ability to provide real-time physiological and behavioral data, interoperability with third-party healthcare applications, and accessibility through APIs or SDKs. Due to space constraints, Table V shows only the selected devices. Specifically, we selected the Withings ScanWatch 2 as the primary wearable device. This choice was driven by its ability to record a single-lead ECG directly from the wrist, enabling early detection of arrhythmias such as atrial fibrillation. This device also provides continuous monitoring, e.g., of heart rate and SpO2, offering a comprehensive cardiovascular assessment. A key advantage of ScanWatch 2 is its integration with IoT platforms via Withings’ well-documented APIs. These APIs ensure secure access to real-time health data (e.g., ECG traces), facilitating efficient data collection and analysis. Moreover, ScanWatch 2 meets stringent medical certification standards, ensuring measurement accuracy and compliance with regulatory requirements. Compared to other smartwatches, ScanWatch 2 offers a balance between advanced medical-grade features and ease of integration, making it an optimal solution for telemonitoring. For similar reasons, we chose the Withings Body+ scale. Further details on the selection process of the scale are not provided for space reasons.
@@
Withings ScanWatch 2 - ECG - Heart Rate - Body Temperature - SpO2 - Physical Activity - Sleep Monitoring Yes 30
Withings Body+ - Weight - Body Mass Index - Fat mass - Muscle mass - Bone mass - Water percentage Yes 540
@@ V-C O3 - Developing tools to support decision-making
We implemented the predictive model for stratification using clinical and echocardiographic features. The used features are reported in Table VI. Since we did not find a public dataset to train such a model, we had to use a dataset of the THE ecosystem. This choice could have affected the external validity of the results because the performances of our model might differ in other populations with varying demographic profiles. The original dataset contained 1040 elements that became 357 after pre-processing.
The model architecture is based on an ensemble learning technique, specifically a variation of stacking [46]. In standard stacking, multiple predictive models (often different from each other) are trained on the same set of features, and a meta-model then combines their outputs to generate the final prediction. Differently, we employed two specialized models trained separately: one using only clinical features and the other using only echocardiographic features. The outputs of these two models serve as inputs for the meta-model (a custom version of a voting classifier), which provides the final stratification of patients.
The performance of our model was assessed using metrics that are commonly employed in the ML and medical domains. Specifically, we considered: (i) Accuracy ($\frac{TP+TN}{TP+TN+FP+FN}$), (ii) Precision ($\frac{TP}{TP+FP}$), (iii) Sensitivity/Recall ($\frac{TP}{TP+FN}$), (iv) F1-Score ($2\times\frac{\text{Precision}\times\text{Sensitivity}}{\text{Precision}+\text{Sensitivity}}$), and (v) Diagnostic Odds Ratio ($\frac{TP\times TN}{FP\times FN}$). For the sake of clarity, TPs (True Positives) refer to patients correctly classified as people at HF risk, while TNs (True Negatives) refer to patients correctly identified as not at HF risk. Instead, FPs (False Positives) represent patients mistakenly classified as at risk, while FNs (False Negatives) represent patients mistakenly identified as not at risk.
@@
Metric Value
Accuracy 78%
Precision 70%
Sensitivity 91%
F1-Score 79%
Diagnostic Odds Ratio 20
The experimental results of our model are summarized in Table VII. These results suggest that the model performs pretty well. Although an accuracy of 78% could be considered moderate in some contexts, it is acceptable in the PrediHealth project given the priority of capturing patient at HF risk. The high sensitivity (91%) ensures that nearly all HF risk patients are detected, while a relatively low precision (70%) could represent an acceptable trade-off in our project. Although participation in the telemonitoring program could be expensive, it is better to include patients mistakenly classified as at risk (FPs) rather than not including patients mistakenly identified as not at risk (FNs) in such a program. The good balance between sensitivity and precision is shown by the obtained F1-Score value (79%). The high Diagnostic Odds Ratio (20) confirms the robustness of the model to discriminate patients in at HF risk or not. Further information on the model is not provided for space constraints and because it is out of the scope of this paper.
As for the thresholds/scores for risk factors for HF exacerbations when using telemonitoring kits, we chose them on clinical evidence and guidelines [47, 48, 49]. Specifically, threshold values considered at risk include SpO2 $<$ 92%, resting heart rate $>$ 100 bpm (tachycardia) or $<$ 50 bpm (bradycardia), and persistently low HRV, typically Standard Deviation of the Normal-to-Normal $<$ 20 ms. Early signs of decompensation can be inferred when body weight increases by more than 2 kg in three days, suggesting fluid overload. Other critical indicators include systolic blood pressure $>$ 140 mmHg or $<$ 90 mmHg, and diastolic pressure $>$ 90 mmHg or $<$ 60 mmHg, which may indicate hemodynamic imbalance. Finally, a respiratory rate $>$ 20 breaths/min and body temperature $>$ 37.5 or $<$ 36 may signal infection, hypoperfusion, or cardiovascular stress. Finally, a single-lead ECG, validated for atrial fibrillation detection by identifying irregular RR intervals and absent P waves, enables early recognition of AF, which is critical as it can impair cardiac output and increase the risk of decompensation and thromboembolic events. When risk factors occur, the IoT platform will notify both the patient and the appropriate healthcare infrastructure. This phase of PrediHealth is still ongoing.
@@ VI Final Remarks and Expected Impact
The PrediHealth project is centered around the development of a web-based IoT application designed for remote Heart Failure (HF) patient monitoring. This innovative solution enables the continuous monitoring of vital signs and environmental parameters, thereby improving patient engagement and allowing for real-time oversight. By leveraging AI/ML-powered algorithms, we provide support to clinical decision-making (e.g., patient stratification), monitor disease, and proactively prevent exacerbations. With its holistic and data-driven approach to integrating clinical and environmental factors, PrediHealth has the potential to transform HF management, foster personalized care, and contribute to long-term healthcare sustainability.
@@ Acknowledgment
This project has been financially supported by the European Union NEXTGenerationEU project and by the Italian Ministry of the University and Research MUR (M4C2) as part of the project “Tuscany Health Ecosystem” - THE - Spoke 10 - CUP - J13C22000420001. One of the authors involved in the presented research is funded by the project “DHEAL – COM- Digital Health Solutions in Community Medicine” under the Innovative Health Ecosystem (PNC) - National Recovery and Resilience Plan (NRRP) program funded by the Italian Ministry of Health.